Initially, a car travels along a straight road with a speed of 35 m/s. if the brakes are applied and the speed of the car is reduced to 10 m/s in 15 s, determine the constant deceleration of the car.
1
Expert's answer
2020-08-24T13:22:04-0400

Initial velocity"(u)"= "35 m\/s"

Final Velocity"(v)=" "10 m\/s"

Time taken"(t)=" "15 s"

Acceleration;


"a=" "\frac{v-u}{t}"


"a=\frac{10-35}{15}=-1.67"

Therefore the constant deceleration of the car is "-1.67m\/s"